OpenStudy (anonymous):

When using synthetic division to do the following problem, what number should be used as the divisor?

OpenStudy (amistre64):

you want to zero out the denominator ... so if the poly you are dividing by is say: (x-c) then x-c = 0 when x=c, use c in the synthetic process

OpenStudy (radar):

Division by 0 is not allowed!

OpenStudy (radar):

Just sayin, as I see no divisor!

OpenStudy (anonymous):

x+4 so it would be -4 for my answer?

OpenStudy (amistre64):

i divide by zero all the time :) -4 yes

OpenStudy (amistre64):

the division process assumes that (x+4) is a factor, and that if (x+4) = 0, then the poly is zero as well

OpenStudy (anonymous):

thank you. you only do that for synthetic division?

OpenStudy (amistre64):

the remainder makes up the difference so its not really a divide by zero application :)

OpenStudy (amistre64):

\[\frac{P(x)}{d(x)}=Q(x)+\frac{r(x)}{d(x)}\] \[d(x)\frac{P(x)}{d(x)}=Q(x)d(x)+d(x)\frac{r(x)}{d(x)}\] \[P(x)=Q(x)d(x)+r(x)\] the limit of d(x) to 0 gives us the r(x)

OpenStudy (amistre64):

its applicable to synthetic division ... im not sure if you do it ONLY for synthD

OpenStudy (radar):

@mocham Why not post the "following problem" as an example so it can be discussed.

OpenStudy (anonymous):

hmmm here are my options: A. 1 B. -1 C. 4 D. -4 E. None of the above -4 is in there ....

OpenStudy (radar):

@mocham your original post states that there is a particular problem , never mind as the question is now closed..